# Dongyue Lu1, Lingdong Kong1, Gim Hee Lee1, Camille Simon Chane2, Wei Tsang Ooi1

### v1

@@ Abstract
Event cameras offer unparalleled advantages for real-time perception in dynamic environments, thanks to their microsecond-level temporal resolution and asynchronous operation. Existing event-based object detection methods, however, are limited by fixed-frequency paradigms and fail to fully exploit the high-temporal resolution and adaptability of event cameras. To address these limitations, we propose FlexEvent, a novel event camera object detection framework that enables detection at arbitrary frequencies. Our approach consists of two key components: FlexFuser, an adaptive event-frame fusion module that integrates high-frequency event data with rich semantic information from RGB frames, and FAL, a frequency-adaptive learning mechanism that generates frequency-adjusted labels to enhance model generalization across varying operational frequencies. This combination allows our method to detect objects with high accuracy in both fast-moving and static scenarios, while adapting to dynamic environments. Extensive experiments on large-scale event camera datasets demonstrate that our approach surpasses state-of-the-art methods, achieving significant improvements in both standard and high-frequency settings. Notably, our method maintains robust performance when scaling from $20$ Hz to $90$ Hz and delivers accurate detection up to $180$ Hz, proving its effectiveness in extreme conditions. Our framework sets a new benchmark for event-based object detection and paves the way for more adaptable, real-time vision systems.
Figure 1: Event camera detection at varying frequencies. The performance of the classic RVT detector (Gehrig & Scaramuzza, 2023) drops significantly at higher event operational frequencies. Motivated by this, we propose FlexEvent, a robust and flexible detector that maintain high accuracy across a wide range of frequencies, ensuring strong adaptability in dynamic sensing environments.
@@ 1 Introduction
Event cameras have garnered significant attention for their ability to capture dynamic scenes with microsecond-level temporal resolution (Gallego et al., 2022). Unlike conventional RGB cameras, which capture entire frames at fixed intervals, event cameras operate asynchronously, responding to changes in pixel intensity at each location (Zou et al., 2022). This low-latency operation reduces motion blur and enables highly energy-efficient sensing, making event cameras ideal for real-time applications such as autonomous driving, robotics, and surveillance (Steffen et al., 2019; Kong et al., 2024b; Chen et al., 2023).
@@
Despite their potential, existing event camera object detection methods often fail to fully leverage the high-frequency temporal information captured by event cameras (Cordone et al., 2022; Gehrig & Scaramuzza, 2022; Jeziorek et al., 2023). Most approaches align event data with the lower frequency of RGB cameras by adopting a fixed time interval between event streams and frame-based annotations (Perot et al., 2020; Gehrig et al., 2021b). While this strategy simplifies data processing, it inevitably overlooks the rich temporal details embedded in high-frequency event streams. Given that human annotations are often synchronized with slower frame rates, current detection models miss valuable information from high-frequency event data, resulting in suboptimal performance when rapid object detection is required in dynamic environments (Messikommer et al., 2020; Schaefer et al., 2022).
@@
To address these limitations, we introduce FlexEvent, a novel event camera object detection framework designed to tackle the challenging problem of object detection at varying operational frequencies. Our approach addresses the need for high-frequency detection in fast-changing environments, while adapting to different operational frequencies. We propose two key innovations: (1) FlexFuser, an adaptive event-frame fusion module, and (2) FAL, a frequency-adaptive learning mechanism.
@@
Flexible Event-Frame Fusion. The first component, FlexFuser, addresses the limitations of event data, which often lacks semantic and texture-rich information, especially at higher frequencies (Zhou et al., 2023), by integrating the rich spatial and semantic information from RGB frames with the high-temporal resolution of event streams. It enables high detection accuracy even in fast-moving environments. Furthermore, training on high-frequency event data is computationally expensive and impractical due to the significant human effort required to label such data. FlexFuser mitigates this by sampling event data at varying frequencies, aligning them with the normal frame rate during training, thus maintaining efficiency while preserving the high-frequency benefits at inference time.
@@
Frequency-Adaptive Learning. The second component, FAL, enhances the generalization capability of event camera detectors across varying operational frequencies, by generating frequency-adjusted labels for the unlabeled high-frequency data. These labels allow the model to learn from high-frequency event streams without manual annotations, and iterative refinement through self-training ensures that the model remains robust across different motion dynamics and frequency settings. Together, these two components allow for accurate real-time detection in rapid scene changes and adapt to a wide range of operational frequencies, by leveraging the temporal richness of event data and the semantic detail of RGB frames.
@@
Our extensive experiments validate the effectiveness of FlexEvent on multiple large-scale event camera datasets. Our approach consistently outperforms recent detectors across both standard and high-frequency settings. In particular, we achieve mAP gains of $\mathbf{15.5\%}$, $\mathbf{9.4\%}$, and $\mathbf{10.3\%}$ over the previous best-performing detectors on the DSEC-Det (Gehrig & Scaramuzza, 2024), DSEC-Detection (Tomy et al., 2022), and DSEC-MOD (Zhou et al., 2023) datasets, respectively. Our model also maintains $\mathbf{96.2\%}$ of its performance when the operational frequency shifts from $\mathbf{20}$ Hz to $\mathbf{90}$ Hz, and delivers accurate detection at frequencies as high as $\mathbf{180}$ Hz, proving its robustness under extreme conditions.
[…]
@@ 2 Related Work
Event Camera Object Detection. Event-based detection methods can be broadly split into two approaches: GNNs/SNNs and dense feed-forward models. GNNs build dynamic spatio-temporal graphs by subsampling events (Gehrig & Scaramuzza, 2022; Sun & Ji, 2023; Messikommer et al., 2020; Schaefer et al., 2022), but they face challenges in propagating information over large spatio-temporal regions, especially for slow-moving objects. SNNs offer efficient sparse information transmission but are often hindered by their non-differentiable nature, complicating optimization processes (Cuadrado et al., 2023; Cordone et al., 2022; Zhang et al., 2022). Dense, feed-forward models represent the second approach. Initial methods using fixed temporal windows (Chen, 2018; Iacono et al., 2018; Jiang et al., 2019) struggle with slow-moving or stationary objects due to their limited capability to capture long-term temporal data. Subsequent advancements incorporate RNNs and transformers to enhance temporal modeling capabilities (Perot et al., 2020; Zubić et al., 2023; Li et al., 2022; Gehrig & Scaramuzza, 2023; Peng et al., 2024), but these models often still lack semantic richness and face difficulties in adapting to variable frequencies and highly dynamic scenarios.
@@
Event-Frame Multimodal Learning. To overcome the limited texture in event streams, multimodal fusion techniques combining event-based and frame-based data have gained traction across tasks, such as deblurring (Sun et al., 2022a; Zhang et al., 2020), depth estimation (Gehrig et al., 2021a; Uddin et al., 2022), and tracking (Zhao et al., 2022; Gehrig et al., 2020). Earlier object detection approaches fuse event and image data during post-processing (Li et al., 2019; Chen et al., 2019), but they lack meaningful feature-level interaction. Recent works focus on deeper feature fusion (Tomy et al., 2022; Cao et al., 2022; 2021), with advanced methods introducing pixel-level spatial attention or temporal transformers for asynchronous processing (Zhou et al., 2023; Li et al., 2023; Gehrig & Scaramuzza, 2024; Cao et al., 2024). Some approaches (Li et al., 2023; Gehrig et al., 2021a) explore combining events and frames through asynchronous multi-modal fusion, enabling inference at varying frequencies. However, they do not focus on high-frequency event data and fully leverage the temporally rich nature of event streams. Furthermore, these methods still face challenges in fully exploiting complementary strengths and addressing feature imbalance in event-frame detection. Unlike previous methods, FlexEvent employs a more comprehensive fusion framework that effectively combines high-temporal resolution event data with rich semantics from RGB frames, enabling robust object detection across varying frequencies while addressing feature imbalance.
@@
Label-Efficient Learning in Event Data. Due to limited annotated datasets, label-efficient learning has become an important area for event-based vision. Several studies attempt to reconstruct images from event data (Rebecq et al., 2019; 2021; Stoffregen et al., 2020) or leverage knowledge distillation from pre-trained frame-based models (Wang et al., 2021; Sun et al., 2022b; Yang et al., 2023; Kong et al., 2024a). Other approaches utilize pre-trained models or self-supervised losses (Klenk et al., 2022; Wu et al., 2023; Zhu et al., 2019). LEOD (Wu et al., 2024) pioneers object detection with limited labels but does not address high-frequency generalization. A recent state-space model (Zubić et al., 2024) adapts to varying frequencies without retraining but struggles to detect static objects at high frequencies due to its reliance solely on event data. In contrast, FlexEvent is specifically designed to adapt to varying event frequencies, ensuring consistent performance even in scenarios with limited labels, and effectively detecting both stationary and fast-moving objects.
3 FlexEvent: A Flexible Event Object Detector
@@
In this section, we elaborate on the technical details of our FlexEvent framework. We start with the foundational concepts of event data and their representation in Sec. 3.1. We then introduce the FlexFuser module in Sec. 3.2, which adaptively fuses event and frame data to enhance detection across varying frequencies. Finally, we detail the frequency-adaptive learning (FAL) mechanism in Sec. 3.3, which enables our model to generalize effectively across diverse temporal conditions using self-training and adaptive label generation. The overall framework is illustrated in Fig.2.
@@
Figure 2: Framework Overview. The proposed FlexEvent consists of two branches: Event and Frame. The event branch captures high-temporal resolution data, while the frame branch leverages the rich semantic information from frames (cf. Sec. 3.1). These branches are fused dynamically through FlexFuser, allowing adaptive integration of event and frame data (cf. Sec. 3.2). Additionally, the frequency-adaptive learning (FAL) mechanism ensures robust detection performance across varying operational frequencies (cf. Sec. 3.3). Together, these components enable the model to handle diverse motion dynamics and maintain high detection accuracy in varying frequency scenarios.
@@ 3.1 Background & Preliminaries
Event Processing. Event cameras are bio-inspired vision sensors that capture changes in log intensity per pixel asynchronously, rather than capturing entire frames at fixed intervals. Formally, let $I(x,y,t)$ denotes the log intensity at pixel coordinates $(x,y)$ and time $t$. An event $e$ is generated at $(x,y,t)$ whenever the change in log intensity $\Delta I$ exceeds a certain threshold $C$. Such a process can be modeled as:
@@
To leverage event data with convolutional neural network layers, we preprocess events into a 4D tensor $E$ with dimensions representing the polarity, temporal discretization $T$, and spatial dimensions $(H,W)$. This representation involves mapping a set of events $\mathcal{E}$ within time interval $[t_{a},t_{b})$ into:
$E(p,\tau,x,y)=\sum_{e_{k}\in\mathcal{E}}\delta(p-p_{k})\delta(x-x_{k},y-y_{k})\delta(\tau-\tau_{k}),$ (2)
@@
where $\tau_{k}=\left\lfloor\frac{t_{k}-t_{a}}{t_{b}-t_{a}}\cdot T\right\rfloor$. This tensor captures event activity in $T$ discrete time slices, yielding a compact representation suitable for 2D convolutions by flattening the polarity and temporal dimensions.
@@
Problem Formulation. Given two consecutive frames $F_{a}$ and $F_{b}$ captured at timestamps $T_{a}$ and $T_{b}$, our objective is to leverage the event stream over the interval $[T_{a},T_{b}]$ to detect objects at the end timestamp $T_{b}$. Existing event-based object detection methods often use fixed time intervals $\Delta T$, limiting adaptability to dynamic environments (Perot et al., 2020). Additionally, integrating semantic information from RGB frames remains challenging, affecting performance in complex scenarios (De Tournemire et al., 2020). To address this, we synchronize the event data with frames and explore varying training frequencies, leveraging the temporal richness of event cameras to improve detection accuracy.
3.2 FlexFuser: An Adaptive Event-Frame Fusion Module
@@
In dynamic environments, object detection systems must adapt to varying motion frequencies (Sun et al., 2022a). While event cameras excel at capturing rapid changes in pixel intensity, they often lack the rich spatial and semantic information provided by frames. To address this limitation and fully leverage the complementary strengths of both modalities, we introduce FlexFuser, an adaptive fusion module designed to dynamically combine event data at different frequencies with frame data.
@@
Dynamic Event Aggregation. Given a dataset $\mathcal{D}$, consisting of sequences of calibrated event camera data and frame data with a resolution of $H\times W$, along with corresponding bounding box annotations $y$ collected at frequency $a$, we begin by selecting a batch of frame data $\mathbf{F}$ paired with event data $\mathbf{E}^{a}$, both captured at frequency $a$. To aggregate event data from a higher frequency $b$ (where $b>a$), we divide the time interval $\Delta T^{a}$ corresponding to $\mathbf{E}^{a}$ into $b/a$ smaller sub-intervals. From each sub-interval, we obtain a high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, as defined in Eq. 2.
From this set, we randomly sample one event data point [^1] For simplicity, we use $\mathbf{E}^{b}$ to represent a sample from the set of high-frequency event data $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, rather than explicitly referencing each individual sample from the event set. The same applies to other frequencies. $\mathbf{E}^{b}$. By doing so, the sampled high-frequency event data $\mathbf{E}^{b}$ is temporally aligned with the frame data $\mathbf{F}$ and the base frequency event data $\mathbf{E}^{a}$. This synchronization of event streams at different frequencies ensures consistent and reliable processing for subsequent stages.
Figure 3: Illustration of the FlexFuser module. We show a general example of event and frame under frequency $a$ at stage $i$.
@@
Feature Extraction. Let $\phi_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}(\cdot)$ and $\phi_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}(\cdot)$ represent the event- and frame-based networks, respectively, where the former employs the RVT (Gehrig & Scaramuzza, 2023) for extracting features from event data, and the latter uses ResNet-50 (He et al., 2016) for feature extraction from frames. Both networks are structured into four stages, as shown in Fig. 2.
At each scale $i$, we extract features ${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a},^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{b}$ from the event data and ${}^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}$ from the frame data:
${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a}=\phi_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{(i)}(\mathbf{E}^{a}),\quad^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{b}=\phi_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{(i)}(\mathbf{E}^{b}),\quad^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}=\phi_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}^{(i)}(\mathbf{F}),$ (3)
@@
where ${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a}$ and ${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{b}\in\mathbb{R}^{B\times C_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}\times H_{i}\times W_{i}}$, ${}^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\in\mathbb{R}^{B\times C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\times H_{i}\times W_{i}}$. Here, $i$ denotes the scale, $B$ is the batch size, and $C_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}$ and $C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}$ are the dimensions of the feature maps extracted from the event and frame data, respectively.
@@
Event-Frame Adaptive Fuser. To effectively fuse the event and frame data, we employ an adaptive fuser that is consistent across different event data frequencies. At each scale $i$, taking the low-frequency event features ${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a}$ as an example, we concatenate the feature maps from both the event and frame branches as follows:

### v3

# FlexEvent: Towards Flexible Event-Frame Object Detection at Varying Operational Frequencies

@@ Abstract
Event cameras offer unparalleled advantages for real-time perception in dynamic environments, thanks to the microsecond-level temporal resolution and asynchronous operation. Existing event detectors, however, are limited by fixed-frequency paradigms and fail to fully exploit the high-temporal resolution and adaptability of event data. To address these limitations, we propose FlexEvent, a novel framework that enables detection at varying frequencies. Our approach consists of two key components: FlexFuse, an adaptive event-frame fusion module that integrates high-frequency event data with rich semantic information from RGB frames, and FlexTune, a frequency-adaptive fine-tuning mechanism that generates frequency-adjusted labels to enhance model generalization across varying operational frequencies. This combination allows our method to detect objects with high accuracy in both fast-moving and static scenarios, while adapting to dynamic environments. Extensive experiments on large-scale event camera datasets demonstrate that our approach surpasses state-of-the-art methods, achieving significant improvements in both standard and high-frequency settings. Notably, our method maintains robust performance when scaling from $20$ Hz to $90$ Hz and delivers accurate detection up to $180$ Hz, proving its effectiveness in extreme conditions. Our framework sets a new benchmark for event-based object detection and paves the way for more adaptable, real-time vision systems.
Figure 1: Illustrative examples of event camera object detection at varying frequencies. The detection performance of the classic RVT detector [18] tends to drop significantly at higher event operational frequencies. Motivated by this observation, we propose FlexEvent, a robust and flexible event-frame detector that maintains high detection accuracy across a wide range of frequencies (low, middle, high), ensuring strong adaptability in real-world, dynamic sensing environments.
@@ 1 Introduction
Event cameras have garnered significant attention for their ability to capture dynamic scenes with microsecond-level temporal resolution [12]. Unlike RGB cameras, which capture entire frames at fixed intervals, event cameras operate asynchronously, responding to changes in pixel intensity at each location [68]. This low-latency operation reduces motion blur and enables highly energy-efficient sensing, making event cameras ideal for real-time applications [45, 30, 6].
@@
Despite their potential, existing event detectors often fail to fully leverage the high-frequency temporal information [8, 15, 23]. Most approaches align event data with the lower frequency of frames by adopting fixed time intervals between event streams and frame-based annotations [40, 17]. While this simplifies data processing, it inevitably overlooks the rich temporal details embedded in high-frequency event streams. Given that human annotations are often synchronized with slower frame rates, current detectors miss valuable information from higher frequencies, resulting in suboptimal performance when rapid object detection is required in dynamic environments [36, 43].
@@
To address these limitations, we introduce FlexEvent, a novel framework designed to tackle the challenging problem of object detection at varying operational frequencies. Our approach addresses the need for high-frequency detection in fast-changing environments, while adapting to different operational frequencies. We propose two key innovations: 1) an adaptive event-frame fusion module, and 2) a frequency-adaptive fine-tuning mechanism.
@@
The first component, FlexFuse, addresses the limitations of event data, which often lacks semantic and texture-rich information, especially at higher frequencies [66], by integrating the rich spatial and semantic information from RGB frames with the high-temporal resolution of event streams. It enables high detection accuracy even in fast-moving environments. Furthermore, training on high-frequency event data is computationally expensive and impractical due to the significant human effort required to label such data. FlexFuse mitigates this by sampling event data at varying frequencies, aligning them with the normal frame rate during training, thus maintaining efficiency while preserving the high-frequency benefits at inference time.
@@
The second component, FlexTune, enhances the generalization capability of event camera detectors across varying operational frequencies, by generating frequency-adjusted labels for the unlabeled high-frequency data. These labels allow the model to learn from high-frequency event streams without manual annotations, and iterative refinement through self-training ensures that the model remains robust across different motion dynamics and frequency settings. Together, these two components allow for accurate real-time detection in rapid scene changes and adapt to a wide range of operational frequencies, by leveraging the temporal richness of event data and the semantic detail of RGB frames.
@@
Our extensive experiments validate the effectiveness of FlexEvent on multiple large-scale event camera datasets. Our approach consistently outperforms recent detectors across both standard and high-frequency settings. In particular, we achieve mAP gains of $\mathbf{15.5\%}$, $\mathbf{9.4\%}$, and $\mathbf{10.3\%}$ over the previous best-performing detectors on DSEC-Det [16], DSEC-Detection [50], and DSEC-MOD [66], respectively. Our model also maintains $\mathbf{96.2\%}$ of its performance when the operational frequency shifts from $\mathbf{20}$ Hz to $\mathbf{90}$ Hz, and delivers accurate detection at frequencies as high as $\mathbf{180}$ Hz, proving its robustness under extreme conditions. In summary, our contributions are listed as follows:
$\bullet$ The proposed FlexEvent framework is aimed at tackling the challenging problem of event camera object detection at varying frequencies, being one of the first works to address this challenging problem explicitly in real-world conditions.
$\bullet$ We propose FlexFuse, an adaptive event-frame fusion module that leverages the strengths of both event and frame modalities, enabling more efficient and accurate event-based object detection in dynamic environments.
$\bullet$ We introduce FlexTune, a frequency-adaptive fine-tuning mechanism that generates frequency-adjusted labels and improves generalization across various motion frequencies.
$\bullet$ We demonstrate that our approach achieves promising results across large-scale datasets, particularly in high-frequency scenarios, validating its effectiveness to handle real-world, safety-critical problems.
@@ 2 Related Work
Event Camera Object Detection. Event-based detectors can be broadly split into two approaches: GNNs/SNNs and dense feed-forward models. GNNs build dynamic spatio-temporal graphs by sub-sampling events [15, 47, 36, 43], but they face challenges in propagating information over large spatio-temporal regions, especially for slow-moving objects. SNNs offer efficient sparse information transmission but are often hindered by their non-differentiable nature, complicating optimization processes [9, 8, 59]. Dense, feed-forward models represent the second approach. Initial methods using fixed temporal windows [5, 22, 25, 19] struggle with slow-moving or stationary objects due to their limited capability to capture long-term temporal data. Subsequent work incorporates RNNs and transformers to enhance temporal modeling [40, 69, 34, 18, 39], but these models still lack semantic richness and face difficulties in adapting to variable frequencies and highly dynamic scenarios.
@@
Event-Frame Multimodal Learning. Combining events with frame-based data has proven effective in tasks such as deblurring [48, 61], depth estimation [14, 51], and tracking [62, 13]. Early fusion methods perform post-processing combinations [33, 4], but lack feature-level interaction. More recent works propose deeper feature fusion techniques [50, 1, 2], introducing pixel-level attention or temporal transformers for asynchronous processing [66, 32, 16, 3]. Some explore asynchronous multi-modal fusion for flexible inference at different frequencies [32, 14, 60], yet they do not explicitly address high-frequency event data or fully exploit temporal richness. Additionally, balancing contributions from event and frame modalities remains challenging. Unlike these works, FlexEvent introduces a unified fusion framework that adaptively integrates high-frequency event data with semantic-rich frames, achieving robust detection across diverse frequencies while addressing feature imbalance.
@@
Label-Efficient Learning from Event Data. Due to limited annotated datasets, label-efficient learning has become an important area for event-based vision. Several studies attempt to reconstruct images from event data [41, 42, 46] or distill knowledge from pre-trained frame-based models [52, 49, 57, 29]. Other approaches use pre-trained models or self-supervised losses [28, 56, 67]. LEOD [55] pioneers object detection with limited labels but does not address high-frequency generalization. A recent state-space model [70] adapts to varying frequencies without retraining but struggles to detect static objects at high frequencies due to its reliance solely on event data. In contrast, FlexEvent is specifically designed to adapt to varying event frequencies, ensuring consistent performance even in scenarios with limited labels, and effectively detecting both stationary and fast-moving objects.
3 FlexEvent: A Flexible Event-Frame Object Detector
@@
In this section, we elaborate on the technical details of our FlexEvent framework. We start with the foundational concepts of event data and their representation in Sec. 3.1. We then introduce the FlexFuse module in Sec. 3.2, which adaptively fuses event and frame data to enhance detection across varying frequencies. Finally, we detail FlexTune, our frequency-adaptive fine-tuning mechanism, in Sec. 3.3, which enables our model to generalize effectively across diverse temporal conditions via adaptive label generation. The overall framework is illustrated in Fig. 2.
@@
Figure 2: Framework Overview. The proposed FlexEvent consists of two branches: Event and Frame. The event branch captures high-temporal resolution data, while the frame branch leverages the rich semantic information from frames (cf. Sec. 3.1). These branches are fused dynamically through FlexFuse, allowing adaptive integration of event and frame data (cf. Sec. 3.2). Additionally, the FlexTune learning mechanism ensures robust detection performance across varying operational frequencies (cf. Sec. 3.3). Together, these components enable the model to handle diverse motion dynamics and maintain high detection accuracy in varying frequency scenarios.
@@ 3.1 Background & Preliminaries
Event Processing. Event cameras are bio-inspired vision sensors that capture changes in log intensity per pixel asynchronously, rather than capturing entire frames at fixed intervals. Formally, let $I(x,y,t)$ denote the log intensity at pixel coordinates $(x,y)$ and time $t$. An event $e$ is generated at $(x,y,t)$ whenever the change in log intensity $\Delta I$ exceeds a certain threshold $C$, which is modeled as:
@@
To leverage event data with convolutions, we pre-process events into a 4D tensor $E$ with dimensions [18] representing the polarity, temporal discretization $T$, and spatial dimensions $(H,W)$, respectively. This representation involves mapping a set of events $\mathcal{E}$ within time interval $t_{1},t_{2})$ into the following:
$E(p,\tau,x,y)=\sum\nolimits_{e_{k}\in\mathcal{E}}\delta(p-p_{k})\delta(x-x_{k},y-y_{k})\delta(\tau-\tau_{k}),$ (2)
@@
where $\tau_{k}=\left\lfloor\frac{t_{k}-t_{1}}{t_{2}-t_{1}}\cdot T\right\rfloor$. Such a 4D tensor captures event activity in $T$ discrete time slices, yielding a compact representation suitable for 2D operators by flattening the polarity and temporal dimensions.
@@
Problem Formulation. Given two consecutive frames $F_{1}$ and $F_{2}$ captured at timestamps $t_{1}$ and $t_{2}$, our objective is to leverage the event stream over the interval $t_{1},t_{2}]$ to detect objects at the end timestamp $t_{2}$. Existing event detectors use fixed time intervals $\Delta t$, limiting adaptability to dynamic environments [40]. Additionally, integrating semantic information from RGB frames remains challenging, affecting performance in complex scenarios [10]. To address this, we propose to synchronize the event data with frames and explore varying training frequencies, leveraging the temporal richness of event cameras to improve detection accuracy.
3.2 FlexFuse: Adaptive Event-Frame Fusion
@@
In dynamic environments, object detection systems must adapt to varying motion frequencies [48]. While event cameras excel at capturing rapid changes in pixel intensity, they often lack rich spatial and semantic information provided by frames. To address this limitation and fully leverage the complementary strengths of both modalities, we introduce FlexFuse, an adaptive fusion module designed to dynamically combine event data at different frequencies with frames.
@@
Dynamic Event Aggregation. Given a dataset $\mathcal{D}$, consisting of sequences of calibrated event data and frame data with a resolution of $H\times W$, along with corresponding bounding box annotations $y$ collected at frequency $a$, we begin by selecting a batch of frame data $\mathbf{F}$ paired with event data $\mathbf{E}^{a}$, both captured at frequency $a$. To aggregate event data from a higher frequency $b$ (where $b>a$), we divide the time interval $\Delta t^{a}$ corresponding to $\mathbf{E}^{a}$ into $b/a$ smaller sub-intervals. From each sub-interval, we obtain a high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, as defined in Eq. 2. From this set, we randomly sample one data point [^1] For simplicity, we use $\mathbf{E}^{b}$ to represent a sample from the set of high-frequency event data $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, rather than explicitly referencing each individual sample from the event set. The same applies to other frequencies. $\mathbf{E}^{b}$. As detections are predicted for the end of $\Delta t^{a}$, this strategy pairs each frame with the preceding high‐frequency events while introducing only millisecond‐scale jitter that acts as implicit temporal augmentation, strengthening robustness to real‐world synchronization noise. Consequently, $(\mathbf{F},\mathbf{E}^{a},\mathbf{E}^{b})$ are effectively paired across modalities and frequencies, and this synchronization of image streams with event streams at different frequencies ensures consistent and reliable processing for all subsequent stages.
@@
Feature Extraction. Let $\phi_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}(\cdot)$ and $\phi_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}(\cdot)$ represent the event- and frame-based networks, respectively, where the former employs the RVT [18] for extracting features from event data, and the latter uses ResNet-50 [20] for feature extraction from frames. Both networks are structured into four stages, as shown in Fig. 2. At each scale $i$, we extract features $^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{a},^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{b}$ from the event data and $^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}$ from the frame data. This process is formulated as:
$^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{a}=^{(i)}\phi_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}(\mathbf{E}^{a}),^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{b}=^{(i)}\phi_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}(\mathbf{E}^{b}),^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}=^{(i)}\phi_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}(\mathbf{F}),$ (3)
@@
where $^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{a}$ and $^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{b}\in\mathbb{R}^{B\times C_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}\times H_{i}\times W_{i}}$, $^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}\in\mathbb{R}^{B\times C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}\times H_{i}\times W_{i}}$. Here, $i$ denotes the scale, $B$ is the batch size, and $C_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}$ and $C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}$ are the dimensions of the event and frame feature maps, respectively.
Figure 3: Illustration of the FlexFuse module. We show a general example of event and frame under frequency $a$ at the $i$-stage.
@@
Event-Frame Adaptive Fusion. To effectively fuse the event and frame data, we employ an adaptive fusion that is consistent across different event data frequencies. At each scale $i$, taking the low-frequency event features $^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{a}$ as an example, we concatenate the feature maps from both the event and frame branches as follows: